# Evaluating and Steering Modality Preferences in Multimodal Large Language Model

### v1

@@ Abstract
Multimodal large language models (MLLMs) have achieved remarkable performance on complex tasks with multimodal context. However, it is still understudied whether they exhibit modality preference when processing multimodal contexts. To study this question, we first build a MC2 benchmark under controlled evidence conflict scenarios to systematically evaluate modality preference, which is the tendency to favor one modality over another when making decisions based on multimodal conflicting evidence. Our extensive evaluation reveals that all 18 tested MLLMs generally demonstrate clear modality bias, and modality preference can be influenced by external interventions. An in-depth analysis reveals that the preference direction can be captured within the latent representations of MLLMs. Built on this, we propose a probing and steering method based on representation engineering to explicitly control modality preference without additional fine-tuning or carefully crafted prompts. Our method effectively amplifies modality preference toward a desired direction and applies to downstream tasks such as hallucination mitigation and multimodal machine translation, yielding promising improvements. footnotetext: The code and data are released at https://github.com/EchoDreamer/Modality-Preference.
@@ 1 Introduction
Multimodal Large Language Models (MLLMs) [3, 41, 43, 50] have emerged as a powerful paradigm for processing and reasoning across heterogeneous data modalities (e.g., text, images, video). Recent advances demonstrate their exceptional capabilities, achieving remarkable performance on complex tasks with multimodal contexts, including autonomous web browsing, graphical user interface understanding, and multimodal dialogue system [16, 18, 40]. Despite impressive performance, fundamental questions remain about their modality preference, particularly whether they solve problems by genuinely integrating cross-modal information or exhibiting modality bias by over-relying on a single dominant modality.
To investigate the modality preferences of MLLMs, a line of research evaluates MLLMs by comparing their performance on unimodal inputs (text-only vs. image-only presentations of the same problem) [12]. However, this approach introduces inherent bias, as it fails to assess the model’s ability to process and integrate truly multimodal inputs where both images and text are simultaneously present. Another line of research analyzes the relative contributions of textual instructions and visual context using Shapley value metric [4, 37, 36]. However, we contend that both parts are indispensable for task resolution, making this framework less adequate for determining true modality preferences across multi-modal contexts.
@@
In this paper, we provide a controlled setup to study the modality preference in MLLMs. As shown in the left panel of Figure 1, we introduce a modality context conflict evaluation setting, where an MLLM is asked to answer a question based on a pair of contrasting evidence. In this way, we can determine the modality preference based on the answer given by MLLMs. Moreover, to dismiss the effect of several confounding factors, including question comprehension, single-modality perception, and the internal knowledge of MLLMs, we utilize widely adopted images and select samples that demonstrate accurate question comprehension and reliable single-modality perception. Building upon this, we introduce a semi-automated annotation framework to construct the refined Modality Context Conflict dataset, MC2, which covers eight recognition-level tasks with 2,000 carefully selected samples. Using MC2, we conduct a comprehensive analysis of modality preference across a diverse set of 18 representative MLLMs. Our study reveals several intriguing findings:
  • All tested MLLMs generally demonstrate clear modality bias, as illustrated in the right penal of Figure 1, and the modality bias is affected by model-intrinsic factors and task-specific characteristics.
  • The direction of modality preference in MLLMs can be modulated through external interventions, such as modality degradation or instructional guidance.
  • Latent space analysis confirms that modality preferences manifest as geometrically separable patterns in MLLMs’ representations, which can be captured by distinct vectors.
@@
Built on these, we propose a modality preference probing and steering method based on representation engineering to explicitly amplify the direction of modality preference without the need for additional fine-tuning or prompt design. Experimental results show that the proposed method leads to notable performance improvements on both hallucination mitigation and multimodal machine translation. Our main contributions are summarized as follows:
  • We introduce a Modality Context Conflict benchmark for the systematic evaluation of modality preference in MLLMs.
  • Our systematic analysis reveals intrinsic modality preferences in MLLMs, and we find that such preferences are steerable and can be identified through latent representations, offering insights into reasoning under multimodal conflicts.
  • Inspired by these findings, we propose a method to steer modality preference via representation-level intervention, enabling controllable preference adjustment on MC2 and improving performance on downstream tasks.
@@
Figure 1: Illustrations of evaluating modality preference. Left: Use multimodal conflict pairs to evaluate modality preference. Right: Quantified results for modality preferences. Vision and Text represent vision and text preference, other represents irrelevant prediction.
@@ 2 Related Work
Multimodal Large language Models.
In recent years, Large Language Models (LLMs) have achieved impressive results across a wide spectrum of tasks [33, 53, 15, 47, 19]. Building upon these, Multimodal Large Language Models (MLLMs) have extended such capabilities, demonstrating strong performance in visual question answering, structured information understanding, and visual navigation [55, 51, 32, 44, 52]. Recent research has turned to systematically evaluating their biases and failure modes [30, 1, 5, 4, 9], aiming to better understand their limitations and guide the development of more reliable models.
Modality Preference in MLLMs. Modality preference—whether MLLMs prioritize one modality over another and to what extent they favor a specific modality when resolving multimodal inputs—can reveal how MLLMs jointly reason over different modalities. Some studies have also observed that MLLMs often favor textual inputs while neglecting visual information, which can result in hallucinations [8, 34, 54]. This highlights the importance of investigating modality preference.
To study the modality preferences of MLLMs, one common approach involves evaluating model performance on unimodal inputs—comparing results when only text or only image inputs are provided for the same task [12]. However, this method inherently overlooks how models handle genuinely multimodal scenarios where both modalities are presented simultaneously, thus limiting its effectiveness in capturing true modality preference. Another research direction assesses the relative influence of textual and visual inputs using Shapley value-based attribution methods [4, 36, 37]. Yet, we argue that both modalities are typically essential for solving multimodal tasks, making this approach insufficient for identifying a model’s inherent modality preference in realistic multimodal contexts. Concurrent work evaluates the model’s ability to detect conflict under scenarios involving conflicting multimodal contexts [45]. However, conflict detection is only one aspect of multimodal reasoning and does not fully reflect a model’s overall ability to jointly reason across modalities.
@@
In this work, in order to investigate their modality preferences, we simulate multimodal reasoning by examining the behavior of MLLMs in response to questions under scenarios involving conflicting multimodal contexts. Compared to prior work, we carefully control for potential confounding factors that may influence modality preference, including question understanding, unimodal content recognition, and interference from the model’s internal knowledge. We systematically evaluate modality preference and design a method to controllably steer it. Finally, we demonstrate the effectiveness of our approach on downstream tasks.
@@
Representation Engineering and Activation Editing. Extensive research has shown that large language models (LLMs) encode interpretable concepts, such as sentiment, truthfulness, and stylistic attributes in activation space [31, 35, 39, 42]. Building on this foundation, representation engineering has proven effective for editing, enhancing, or suppressing specific behaviors in LLMs [14, 38, 46, 48, 56]. In this work, we extend this paradigm to a novel setting: controlling modality preference in multimodal large language models (MLLMs). Instead of focusing on abstract properties, our method identifies and manipulates activation directions that are sensitive to modality preference, enabling flexible and targeted control over multimodal reasoning behavior.
@@ 3 The MC2 Benchmark
In this section, we introduce the design and methodology behind the construction of the Multimodal Context Conflict dataset, MC2, intended for evaluating modality preference. We outline the data design philosophy in Section 3.1, followed by the data construction pipeline in Section 3.2 and the evaluation method in Section 3.3.
@@ 3.1 Data Design Philosophy
To evaluate modality context preference, we construct modality context conflict pairs that simulate how contexts from different modalities influence model decisions, thereby revealing modality preference. However, the evaluation process may be influenced by various factors such as the internal knowledge within MLLMs, question comprehension and the ability to understand single-modality context. Therefore, on the one hand, we focus on evaluating recognition-level tasks that require minimal external knowledge, reducing the impact of internal knowledge within MLLMs. On the other hand, we ensure that the evaluated models can accurately answer questions based solely on the context of a single modality, minimizing the influence of confounding factors.
Definitions and Notations: The following definitions and notations are used throughout this paper. In this paper, the image is selected as the vision context,$c^{v}$. We utilize advanced LLMs to generate textual contexts $c^{t}$ that conflict with the vision context $c^{v}$ in relation to the question $q$. $A^{v}$ and $A^{t}$ are answers based on the vision context or textual context.
@@ 3.2 Semi-automated Data Construction Pipeline
In this section, we introduce our semi-automated data construction pipeline, which follows a meticulous and iterative process to ensure the robustness and reliability of the dataset, in line with the design principles outlined in Section 3.1. The dataset is derived from the TDIUC [22] dataset, sourced from MS-COCO [26], widely adopted in model development to ensure the evaluated models can recognize the images. We extract image caption $caption$, vision context $c^{v}$, question $q$, and answer $A^{v}$ for each sample from TDIUC as the foundation for data annotation. The pipeline follows these steps:
Candidate Textual Context Generation. We use DeepSeekV3 [28] and ChatGPT4o-mini [20] to generate distractor answers $A^{t}$ and candidate textual contexts $c^{t}$ supporting the distractor answers. We aim to construct textual contexts that are aligned with the visual context in terms of overall scene semantics, while introducing content-level conflicts with the visual context specifically in relation to the given question. Therefore, the following task template is employed to generate the distractor answers and textual contexts:
USER: <caption> <question> <answer based on vision context>. Generate the distractor answer different from the given answer based on the vision context and generate a textual context to support the distractor answer. Ensure that the scene described in the textual context is consistent with the caption.
ASSISTANT: <output>
For each model, we generate two pairs of $A^{t}$ and $c^{t}$ to facilitate downstream data selection and construction. Detailed prompt templates for different LLMs are provided in Appendix B.1.
Context-based Sample Selecting. We use several basic judge MLLMs, such as LLaVA1.5-7B [29] and QwenVL-7B [6], to select samples. This ensures that all subsequent MLLMs to be evaluated demonstrate high recognition capabilities for both vision and textual contexts. The following task template is used to prompt the judge MLLMs for selecting samples:
USER: <vision/textual context> <question>. Answer question based on the context.
ASSISTANT: <output>
We select only those samples for which all judge MLLMs can correctly answer the given question based on either the vision context or the textual context
@@
Human Verification. Although the previous steps generally yield reliable results, we incorporate manual inspection to ensure the high quality of the data annotation. In this step, we verify the existence of conflicts between $c^{v}$ and $c^{t}$ and ensure that both contexts can correctly direct $q$ to the corresponding answers, $A^{v}$ and $A^{t}$. When errors are found in a sample, we either correct or discard the sample entirely. Each category is cross-verified by three annotators to ensure the reliability of the results. The detailed manual annotation tasks and implementation procedures are provided in Appendix  B.1.
@@
Iterative Refinement. The dataset undergoes multiple rounds of refinement through a feedback loop. This process ensures that the conflicts between different modality contexts are preserved and consistently represented, and MLLMs can correctly understand both the vision and textual contexts. The iterative process helps identify and rectify potential errors, thereby enhancing the dataset quality.
To this end, the dataset is organized into eight categories: positional reasoning, counting and color, attribute recognition, sentiment analysis, activity recognition, sports recognition, and object recognition, with 250 samples in each category. Details of the data annotation format along with sample cases and dataset statistics is presented in Appendix B.3.
3.3 Question Design and Preference evaluation
Question design.
We reformulate the original questions using GPT-4o-mini [21] into binary-choice format. Additionally, to mitigate potential position bias in model outputs, we implement a consistency evaluation strategy. Specifically, for each binary-choice question, we create two versions by swapping the order of the answer choices such as, “How many people are there? A. four. B. five.” and “How many people are there? A. five. B. four.” A model’s response is considered consistent only if it selects different answers in the two versions (i.e., A then B, or B then A). If the model selects the same option in both versions, the response is marked as invalid, indicating possible position bias.
Evaluation for modality preference.
To quantify modality preference, we evaluate whether the model’s prediction is based on visual or textual context according to the following three types:
  • Vision: The prediction aligns with visual context, indicating a preference for vision modality.
  • Text: The prediction aligns with textual context, indicating a preference for text modality.
  • Others: The model gives inconsistent responses or outputs irrelevant answers. These instances are discarded from further analysis.
We use the Vision Ratio to quantify the model’s preference toward the vision modality, defined as: ${S_{\text{vision}}}/{S_{\text{vision}}+S_{\text{text}}}$. Vision Ratio greater than 0.5 indicates that the model tends to favor visual information over text.
LLaMAVision 31.2/52.4 20.4/69.6 2.0/93.2 21.2/66.8 4.0/93.2 35.2/47.2 10.0/82.4 38.8/42.8 20.4/68.4
Owl3-7B 60.8/31.6 16.4/72.4 10.8/85.6 22.0/69.6 8.4/88.0 28.4/60.4 17.2/71.2 60.0/29.6 28.0/63.5
LLaVA-next-7B 10.8/74.4 5.2/85.2 0.8/93.2 3.6/79.6 0.4/90.8 6.0/76.0 4.8/73.6 26.0/46.8 7.2/77.5
Table 1: Results for modality preference. Values are reported as vision/text results for each model.
4 What Modality Preference Has MLLMs Developed?
We use the proposed MC2 dataset to evaluate a wide range of MLLMs, including the proprietary GPT-4o-mini [21] and seventeen open-source MLLMs with varying parameter sizes. For detailed model evaluation, please refer to Appendix C.

### v2

@@ Abstract
Multimodal large language models (MLLMs) have achieved remarkable success on complex multimodal tasks. However, it remains insufficiently explored whether they exhibit modality preference, a tendency to favor one modality over another when processing multimodal contexts. To study this question, we introduce MC2 benchmark, which constructs controlled evidence-conflict scenarios to systematically evaluate modality preference in decision-making. Extensive experiments reveal that all 20 tested MLLMs generally demonstrate clear modality preferences, and such preferences can serve as a useful indicator of downstream task performances of MLLMs. Further analysis shows that modality preference can be controlled by instruction guidance and captured within the latent representations of MLLMs. Built on these insights, we propose a probing and steering method based on representation engineering to explicitly control modality preference without requiring additional fine-tuning. This method effectively amplifies modality preference toward a desired direction and demonstrates promising improvements across multiple downstream applications, including multimodal visual understanding and multimodal machine translation. footnotetext: The code and data are released at https://github.com/EchoDreamer/Modality-Preference.
@@ 1 Introduction
Multimodal Large Language Models (MLLMs; Achiam et al., 2023; Team et al., 2023; Wang et al., 2024; Yin et al., 2024) have emerged as a powerful paradigm for processing and reasoning across heterogeneous data modalities (e.g., text, images, video). Recent advances demonstrate their exceptional capabilities on complex tasks with multimodal contexts, including autonomous web browsing (He et al., 2024), graphical user interface understanding (Hong et al., 2024b), and multimodal dialogue systems (Sun et al., 2022). Despite impressive performance, fundamental questions remain about their modality preference—whether MLLMs tend to rely more heavily on one modality than others, and to what extent they favor a specific modality when resolving multimodal inputs.
To investigate this, one line of work (Fu et al., 2024; Amara et al., 2024) compares model performance on unimodal input, providing either only text or only image input for the same question. Another line of research analyzes the relative contributions of textual and visual context, typically by removing one modality to observe the changes of the downstream performance (Park et al., 2025) or Shapley value (Alishahi et al., 2019; Parcalabescu & Frank, 2024; 2022). However, such settings inherently introduce bias, as they isolate modalities, thus failing to reflect how models process inputs in realistic multimodal scenarios, where information from different modalities naturally co-occur.
@@
In this paper, we provide a controllable setup to study the modality preference in MLLMs. As shown in the left panel of Figure 1, we introduce a modality context conflict setting, where MLLMs are asked to answer a question based on a pair of contrasting evidence from different modalities. In this way, we can determine the modality preference based on the answer given by MLLMs. To dismiss the effect of confounding factors including question comprehension, single-modality perception, and the internal knowledge of MLLMs, we annotate and select perception-level tasks that demonstrate accurate question comprehension and reliable single-modality recognition. Building upon this, we introduce a semi-automated annotation framework to construct a refined Modality Context Conflict dataset, MC2, which covers eight perception-level tasks with 2,000 carefully selected samples. Using MC2, we conduct a comprehensive analysis of modality preference across a diverse set of 20 representative MLLMs. Our study reveals several intriguing findings:
  • Most MLLMs (except Qwen2.5-VL and InternVL3) display text preference, as shown in the right panel of Figure 1, and modality preference can serve as a useful indicator of downstream task performances.
  • Internal attention patterns toward specific modalities give rise to modality preference, and the underlying factors can be traced to the training data recipe and the scale of the MLLMs.
  • Modality preference can be modulated through explicit instruction guidance, and the direction of preference can be captured as geometrically separable patterns in the latent space.
@@
Built on these, we propose a modality preference probing and steering method based on representation engineering (Zou et al., 2023) to explicitly amplify the modality preference without additional fine-tuning. Experimental results show that the proposed method leads to notable performance improvements on multimodal visual understanding and multimodal machine translation. Our main contributions are summarized as follows:
@@
Figure 1: Illustrations of evaluating modality preference. Left: Using multimodal conflict context pairs to evaluate modality preference. Right: Quantified scores for Vision and Text modalities, where a higher score indicates a stronger preference toward the corresponding modality. Other represents irrelevant predictions, which are discarded during the statistics.
  • We introduce MC2 to comprehensively evaluate modality preferences in MLLMs and highlight the significance of modality preference in correlating the downstream task performance.
  • Our analysis reveals that intrinsic modality preferences in MLLMs are steerable and identifiable through latent representation, providing insights into multimodal reasoning.
  • We propose a training-free method that steers modality preference via representation-level intervention, enabling controllable preference adjustment and enhancing performance on downstream tasks.
@@ 2 Related Work
2.1 Modality Preference
Existing studies on the modality preference of MLLMs can be broadly divided into two categories: 1) investigating the data-related factors that give rise to modality preference or bias, and 2) analyzing the intrinsic characteristics of modality preference within models.
Data factors influencing modality preference. Research on data-related factors (Guo et al., 2023; Chen et al., 2024; Leng et al., 2024) explores how properties of multimodal datasets give rise to and reinforce modality preference. In particular, many samples in multimodal datasets can be resolved correctly by relying on information from only a single modality. When prevalent in training data, such samples bias the optimization dynamics, encouraging models to disproportionately rely on a single modality (Chen et al., 2024). Furthermore, their inclusion in evaluation benchmarks artificially inflates performance metrics, as models can exploit these unimodal shortcuts instead of performing genuine cross-modal integration (Leng et al., 2024; Winterbottom et al., 2020). While these studies establish data’s role in inducing preference, our work focuses on exploring the intrinsic modality preference inherent in MLLMs themselves, independent of specific data distributions.
Evaluating the intrinsic modality preference in MLLMs. Early studies (Peng et al., 2022; Yang et al., 2024; Wei et al., 2024) analyze modality preference or bias by examining how multimodal models optimize for multimodal inputs. Through such analyses, researchers observe that modality bias has a significant impact on both model optimization and downstream task performance (Peng et al., 2022; Ren et al., 2022; Zhang et al., 2024). While these works offer valuable insights, they typically require training models from scratch, which makes them impractical for large-scale multimodal systems. Recent studies have investigated intrinsic modality preference in MLLMs by evaluating model performance on unimodal inputs—using only text or only image for the same task (Fu et al., 2024; Amara et al., 2024)—and by applying Shapley value-based attribution methods to quantify the contribution of each modality (Alishahi et al., 2019; Parcalabescu & Frank, 2022; 2024). However, in real-world multimodal applications, all modalities are indispensable for task resolution, making these frameworks inadequate for determining truly modality preference. Wu et al. (2025) evaluate the model’s ability to detect conflict under scenarios involving conflicting multimodal contexts. However, conflict detection is only one facet of multimodal reasoning and does not comprehensively reflect a model’s modality preference when processing multimodal contexts.
@@ 2.1 Modality Preference
In this work, we simulate multimodal reasoning by examining the behavior of MLLMs in response to questions under scenarios involving conflicting multimodal contexts. Compared to prior work, we carefully control confounding variables such as input quality, question-understanding ability, and internal model knowledge, and construct a modality context conflict dataset, enabling a more rigorous evaluation of modality preference and uncovering new insights. Furthermore, we design a flexible method which can controllably steer the modality preference and demonstrate effectiveness across multiple downstream tasks.
2.2 Representation Engineering
@@
Extensive research has shown that large language models (LLMs) encode interpretable concepts, such as sentiment, truthfulness, and stylistic attributes in representation space in LLMs (Liu et al., 2023; Panickssery et al., 2023; Subramani et al., 2022; Turner et al., 2023). Building on this foundation, representation engineering has proven effective for editing, enhancing, or suppressing specific behaviors in LLMs (Greenblatt et al., 2023; Stolfo et al., 2024; Wu et al., 2024; Xu et al., 2024; Zou et al., 2023). In this work, we extend this paradigm to a novel setting: controlling modality preference in multimodal large language models (MLLMs). Instead of focusing on abstract properties, our method identifies and manipulates representation directions that are sensitive to modality preference, enabling flexible and targeted control over multimodal reasoning behavior.
@@ 3 The MC2 Benchmark
In this section, we introduce the design and methodology behind the construction of the Multimodal Context Conflict dataset, MC2, intended for evaluating modality preference. We outline the data design philosophy in Section 3.1, followed by the data construction pipeline in Section 3.2 and the question design and evaluation metric in Section 3.3.
@@ 3.1 Data Design Philosophy
To reveal the modality preference of MLLMs when processing truly multimodal inputs, we construct modality context conflict pairs and carefully control various confounding factors that may affect the measurement of modality preference, such as the internal knowledge of MLLMs, question comprehension, and the understanding of unimodal contexts. Therefore, we focus on evaluating perception-level tasks that require minimal external knowledge, reducing the impact of internal knowledge within MLLMs. Besides, we ensure that the evaluated models can accurately answer questions based solely on the context from a single modality, minimizing the influence of confounding factors.
@@ 3.2 Semi-automated Data Construction Pipeline
In this section, we introduce our semi-automated data construction pipeline, which follows a meticulous and iterative process to ensure the robustness and reliability of the dataset, in line with the design philosophy outlined in Section 3.1. The dataset is derived from the TDIUC (Kafle & Kanan, 2017) dataset, sourced from MS-COCO (Lin et al., 2014), widely adopted in model development to ensure the evaluated models can recognize the images. We select the image as vision context $c^{v}$, question $q$, and answer $A^{v}$ based on the vision context and the image caption capcap for each sample from TDIUC as the foundation for data annotation. The pipeline follows these steps:
Textual Context Construction. Given a sample including $c^{v}$, $q$, $A^{v}$ and capcap, we construct candidate contrastive textual contexts $c^{t}$ that conflict with $c^{v}$ specifically in relation to $q$ but are aligned with the $c^{v}$ and capcap in terms of overall scene semantics. We prompt DeepSeekV3 (Liu et al., 2024a) and ChatGPT4o-mini (Hurst et al., 2024b) to generate a distractor answer $A^{t}$ to $q$, together with $c^{t}$ that plausibly supports $A^{t}$, using carefully crafted instructions.
For each model, we generate two pairs of $A^{t}$ and $c^{t}$ to facilitate downstream data selection. To ensure that all evaluated MLLMs demonstrate strong recognition capabilities for both visual and textual contexts, we employ several basic MLLMs, such as LLaVA1.5-7B (Liu et al., 2024b) and QwenVL-7B (Bai et al., 2023), as judges to select samples that can be correctly understood with respect to $c^{v}$ and $c^{t}$.
@@
Human Verification. We incorporate manual inspection to ensure the high quality of the data annotation. Specifically, we verify the existence of conflicts between $c^{v}$ and $c^{t}$ and ensure that both contexts can correctly direct $q$ to the corresponding answers, $A^{v}$ and $A^{t}$. Each sample is cross-verified by three human annotators to ensure the reliability of the results, and when errors are found, annotators either correct or discard the sample entirely.
@@
Iterative Refinement. The dataset undergoes multiple rounds of refinement through a feedback loop between textual context generation and human verification, which helps identify and rectify potential errors, thereby enhancing the dataset quality.
Modality Context Conflict Dataset. To this end, we construct MC2, a modality context conflict dataset including 2000 samples.
The detailed instruction templates for textual context generation, the detailed manual annotation procedures, the data annotation format along with sample cases and dataset statistics are provided in Appendix B.
3.3 Question Design and Evaluation Metric
Question Design. We reformulate the original questions with ChatGPT-4o-mini (Hurst et al., 2024a) into a binary-choice format. To further reduce potential position bias in multimodal inputs, we adopt a consistent evaluation strategy, similar to Liu et al. (2024d). Concretely, for each question, we construct two versions by swapping the order of the answer choices. A model’s prediction is regarded as consistent only if it selects the same answer in both versions for a sample; otherwise, it is labeled as inconsistent. Such inconsistent samples are discarded from the subsequent measurement of modality preference.
Evaluation Metric. Inspired by prior work on evaluating stylistic or knowledge-related preferences of LLMs and MLLMs through conflict-pair contexts (Li et al., 2024b; Xie et al., 2023; Liu et al., 2025), we extend this idea to evaluate the modality preference by designing a metric that captures how MLLMs respond to conflicting signals from different modalities. More importantly, through the careful design of our benchmark, we establish as a basis that the model can reliably understand both modalities in isolation. As shown in Table 9 and Table 10 in Appendix, all models achieve over 95% accuracy when provided with either textual or visual context.
Building on this, our metric evaluates modality preference by assessing how the model’s responses align with textual or visual input when the two provide conflicting signals. The model’s response is then categorized based on which modality it aligns with: 1) Vision: the response aligns with the visual context; 2) Text: the response aligns with the textual context; 3) Others: the responses are ambiguous, uncertain, or inconsistent with either modality, which are discarded from further analysis. Then, we naturally define the Vision Ratio to quantify the model’s preference toward the vision modality, defined as: ${S_{\text{vision}}}/({S_{\text{vision}}+S_{\text{text}}})$, where $S_{\text{vision}}$ or $S_{\text{text}}$ denotes the score of the vision or text modality, computed as the proportion of samples whose responses are categorized as Vision or Text across the dataset. $S_{others}$ is the proportion of samples whose responses are categorized as Others. Vision Ratio greater than 0.5 indicates that the model tends to favor visual context over text.
4 Modality Preferences in MLLMs